OpenStudy (anonymous):

Can somebody please help me evaluate this indefinite integral? (x^4 -x^3)/sqrt(x)

OpenStudy (aravindg):

Split the fraction.

OpenStudy (anonymous):

probably easiest just to simplify the integral by dividing each term individually by the x^(-1/2)

OpenStudy (anonymous):

err, multiply if it's a negative exponent.

OpenStudy (anonymous):

ok, let me try that, i'll just integrate each term individually over the sqrt(x).

OpenStudy (aravindg):

Good.

OpenStudy (anonymous):

ok, the fractional exponents tripped me up a bit, but this is what I came up with: [8x^(9/2) - 10x^(7/2)]/20 I can show you my steps, if you'd like. Can you guys tell me if this is correct?

OpenStudy (anonymous):

Not quite what I get. I get different coefficients.

OpenStudy (aravindg):

Same here. The exponents are right.

OpenStudy (anonymous):

hmm, well what did you get for coefficients? I had 2x^(9/2)/5 - 2x^(7/2)/4 and I multiplied to get 20 as a common denominator, but I think that's right. Let me check my work over...

OpenStudy (anonymous):

Remember, add one to the exponent, THEN divide by it, for the power rule.

OpenStudy (anonymous):

i'm still confused, if the exponents are -1/2 so +1 would be 1/2 and divided by 1/2 would be 2x^(1/2)/1 right?

OpenStudy (anonymous):

\[\frac{x^4 - x^3}{x^{1/2}} = \frac{x^4}{x^{1/2}}-\frac{x^3}{x^{1/2}} = x^{\frac{7}{2}} - x^{\frac{5}{2}}\]

OpenStudy (anonymous):

That looks strange to me, but the exponents before integration should be 7/2 and 5/2 (4- 1/2, and 3 - 1/2)

OpenStudy (anonymous):

ok so after integration it would be 2x^(9/2)/9 - 2x^(7/2)/7 ?

OpenStudy (anonymous):

14x^(9/2) - 18x^(7/2)/63 ?

OpenStudy (anonymous):

Looks good. I normally leave the coefficients as individual fractions, but your teacher may not want that.

OpenStudy (anonymous):

ok thank you so much for your help! I understand this much more than when we started, so yay!

OpenStudy (anonymous):

Outstanding :) Keep at 'em. They get easier to do :)